OpenStudy (anonymous):

I have a question! Which correctly gives the expression using addition? 14.3 – (–21.2) A. 14.3 + (–21.2) B. 14.3 + 21.2 C. –14.3 + (–21.2) D. –14.3 + 21.2

OpenStudy (anonymous):

b

OpenStudy (anonymous):

how

OpenStudy (anonymous):

@matthew120

OpenStudy (matthew120):

oh hold on

OpenStudy (anonymous):

because both numbers are positive and not negative or they ae not to negatives so therefore its correct

OpenStudy (anonymous):

ill give medal anyone

OpenStudy (matthew120):

b

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

Can I ask a few more?

OpenStudy (anonymous):

What is –26 – (–50.8)?

OpenStudy (anonymous):

fine ill do it myself

OpenStudy (jango_in_dtown):

@MyPhilHillName -26-(-50.8)=-26+50.8=24.8